OpenStudy (halfdeafdarling):

Question super simple but confused me >.< (–5 + 3i) ⋅(1 –2i) <-- i know distributive property is used but everytime I do this i keep getting like -11+13i

OpenStudy (photon336):

You would foil it out like you would any problem like that but you must also use the idea below: well, we know that, \[\sqrt{-1} = i \] so if we square both sides. \[(\sqrt{-1)})^{2} = i^2 \] \[-1 = i^{2}\]

OpenStudy (photon336):